# Filtered Direct Preference Optimization

### v2

@@ 1 Introduction
RLHF operates by taking a preference dataset and a language model (LM) as inputs to produce an LM refined by these preferences (Ouyang et al., 2022). It is broadly divided into two approaches concerning the use of a reward model (RM): RM-based RLHF, which learns an RM from the preference dataset and then uses it to optimize an LM through reinforcement learning (RL), and an RM-free approach that directly adjusts an LM based on preference data. This division mirrors the distinction between offline model-based and model-free RL (Sutton & Barto, 2018). [^1] RM-based RLHF first estimates the environment (specifically, the reward model; we do not need to estimate a state transition model because it is typically predefined in NLG tasks) and then optimizes an LM under the estimated environment. This approach is in itself a form of model-based RL. Each approach offers unique advantages and requires careful application based on specific goals and contexts. For instance, in scenarios with limited data, model-based RL might be preferable due to its data efficiency, though its computational cost is generally higher than that of model-free RL (Moerland et al., 2022; Levine et al., 2020). Consequently, RM-based RLHF may be more effective in leveraging data than RM-free methods.
In this paper, we explore the impact of preference dataset quality on the performance of LMs optimized by DPO, specifically focusing on the quality of response texts rather than labeling accuracy. We demonstrate that DPO is more affected by text quality variations within the dataset than typical RLHF methods, as shown in Figure 1 (A). Notably, we observe that lower-quality data can create performance bottlenecks. In response to this challenge, we introduce a novel approach named filtered direct preference optimization (fDPO), which aims to harness potential data efficiency advantages of RM-based RLHF. It uses a trained RM to identify and discard samples of lower quality than those generated by an LM during fine-tuning. Our experiments, conducted with 160M and 1.4B LMs, show that fDPO significantly enhances the effectiveness of DPO, as illustrated in Figure 1 (B).
@@
Figure 1: Performance comparison of alignment methods using a 160M LM with the AlpacaFarm dataset (Dubois et al., 2023). (A) shows the impact of dataset quality on RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), with DPO exhibiting greater sensitivity to dataset quality variations. The normalized gold rewards are adjusted to scale the initial average reward to 0 and the post-optimization average reward to 1 for clarity. The initial model’s average gold reward was $-2.95$, while the average rewards for RLHF and DPO were $-1.49$ and $-2.31$, respectively. (B) compares the performance of DPO and the proposed fDPO on a mixed-quality dataset, illustrating that fDPO effectively mitigates the impact of data quality variations.
@@
  • Our experiments on instruction following tasks demonstrate that fDPO substantially enhances the performance of LMs.
The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 delves into the background. In Section 4, we detail the proposed method, fDPO, explaining its mechanisms and the rationale behind its design. Section 5 presents the experimental results, illustrating the effectiveness of fDPO and its impact on LM performance. Finally, Section 6 concludes the paper, summarizing our findings, discussing limitations, and suggesting directions for future research.
@@ 2 Related work
We examine methods for aligning LMs with human preferences, focusing on RLHF and its alternatives. Most RLHF approaches utilize an RM (Ouyang et al., 2022; Touvron et al., 2023; Dubois et al., 2023; Casper et al., 2023). These methods fine-tune LMs using RL algorithms such as REINFORCE (Williams, 1992; Rennie et al., 2017), proximal policy optimization (PPO) (Schulman et al., 2017), or their variants Sutton & Barto (2018). However, there are notable reinforcement-learning-free approaches (Zhao et al., 2023; Liu et al., 2024), and learning-free methods that leverage the RM at decoding time, with best-of-N (BoN) sampling being a prominent example (Stiennon et al., 2020; Nakano et al., 2021).
@@
DPO and its generalizations (Azar et al., 2023; Tang et al., 2024) represent significant RM-free methods. Some DPO variants explore different regularizations (Wang et al., 2024) or use a divided dataset for stepwise training (Gou & Nguyen, 2024). Other variants propose adding an offset to the DPO objective function (Amini et al., 2024) or a curriculum learning (Gou & Nguyen, 2024), by evaluating the quality difference between chosen and rejected responses. This focus on response quality is relevant to our approach.
Despite various advancements in DPO, our study is the first, to our knowledge, to explore its significant dependence on preference dataset quality and to attempt to refine the dataset for better performance. Additionally, our proposed fDPO method complements these developments. Integrating fDPO with these methods is an exciting possibility for future work, potentially leading to even more effective ways to align LMs with human preferences.
@@ 3 Background
This section delves into RLHF in Section 3.1 and explores DPO in Section 3.2.
@@ 3.1 Reinforcement Learning from Human Feedback
Reinforcement learning from human feedback (RLHF) frames the application of human feedback to enhance performance of a language model (LM) within the context of an RL problem. The process incorporates a pre-trained LM $\pi_{\theta}(y\,|\,x)$, with $\theta$ denoting model parameters, $x$ the prompt, and $y$ the associated response. It also includes a demonstration dataset $\mathcal{D}_{\rm demo}$ for initial supervised fine-tuning and a preference dataset $\mathcal{D}$ for further RL fine-tuning. The aim is to refine the LM $\pi_{\theta}$ with these datasets $\mathcal{D}_{\rm demo}$ and $\mathcal{D}$. We will present an overview of the widely studied RLHF pipeline (Ouyang et al., 2022), establishing the notations and concepts for understanding our contributions. The RLHF pipeline comprises three principal phases: (1) supervised fine-tuning, (2) reward modeling, and (3) RL fine-tuning.
@@
To estimate the probability that a given response is preferred over another, the RM $r_{\phi}$ utilizes the Bradley-Terry model (Bradley & Terry, 1952), which is formulated as:
$p_{\textrm{BT}}(y_{c}\succ y_{r}\,|\,x,r_{\phi})$ $=\frac{\exp(r_{\phi}(x,y_{c}))}{\exp(r_{\phi}(x,y_{c}))+\exp(r_{\phi}(x,y_{r}))}=\sigma(r_{\phi}(x,y_{c})-r_{\phi}(x,y_{r})),$
where $\sigma(x)=\frac{1}{1+\exp(-x)}$ is the sigmoid function. The RM is trained by minimizing the following negative log-likelihood of the observed preferences in the dataset:
$L(\phi)=-\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}[\log\sigma(r_{\phi}(x,y_{c})-r_{\phi}(x,y_{r}))]$ (1)
@@
The RL fine-tuning phase uses the learned RM $r_{\phi}$ to optimize the SFT model $\pi_{\theta}$. The goal is to enhance $\pi_{\theta}$ by maximizing the expected reward while maintaining closeness to the reference LM $\pi_{\rm ref}$, striking a balance that avoids large deviations from the pre-trained behavior. The SFT model before RL fine-tuning is often used as $\pi_{\rm ref}$. This is achieved through policy gradient methods like proximal policy optimization (PPO) (Schulman et al., 2017). The optimization objective is formalized as
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\bigg{[}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\,|\,x)}[r_{\phi}(x,y)]-\beta D_{\textrm{KL}}(\pi_{\theta}(\cdot\,|\,x),\pi_{\rm ref}(\cdot\,|\,x))\bigg{]},$ (2)
@@
$D_{\textrm{KL}}(p,q)=\mathbb{E}_{x\sim p}\left[\log\frac{p(x)}{q(x)}\right].$
@@ 3.2 Direct Preference Optimization
$L_{\rm DPO}(\theta)=\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\left[\log\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}-\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}\bigg{)}\right],$ (3)
@@ 4 Filtered Direct Preference Optimization
In this section, we propose an approach called filtered direct preference optimization (fDPO), which refines the dataset used in DPO. The principle of fDPO is straightforward: it aims to discard lower-quality samples compared to those generated by the LM, as determined by an proxy RM. This strategy is intuitively derived from observing that lower-quality data can create performance bottlenecks in DPO. First, we give an implementation of fDPO in Section 4.1. Then, we will elaborate on the motivation of fDPO by analyzing DPO’s behavior in Section 4.2.
@@
fDPO needs to assess the quality of responses. For this purpose, a straightforward approach is to use an RM. This incorporation of an RM diverges from the RM-free nature of the original DPO, aligning fDPO closer to RM-based RLHF approaches and making DPO more effective in leveraging data.
@@
1:Pre-trained LM $\pi_{\theta}$, RM $r_{\phi}$, demonstration dataset $\mathcal{D}_{\rm demo}$, preference dataset $\mathcal{D}_{\rm pref}$, maximum epoch $M$, and maximum filtered rate $\gamma$.
@@
6:while $m<M$ and $|\mathcal{D}_{\rm f}|/|\mathcal{D}_{\rm pref}|>1-\gamma$ do
@@ 4.2 Background and Motivation for fDPO
The motivation for developing fDPO stems from the observation that the quality of data in DPO significantly affects the performance of the resulting LM. More specifically, upon differentiating the objective function of DPO in Eq. 3, we obtain
$\nabla_{\theta}L_{DPO}(\theta)=-\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\bigg{[}\underbrace{w_{\theta}(x,y_{c},y_{r})\nabla_{\theta}\log\pi_{\theta}(y_{c}|x)}_{\textrm{increase likelihood of }y_{c}}-\underbrace{w_{\theta}(x,y_{c},y_{r})\nabla_{\theta}\log\pi_{\theta}(y_{r}|x)}_{\textrm{decrease likelihood of }y_{r}}\bigg{]},$
where $w_{\theta}(x,y_{c},y_{r})=\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}-\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}\bigg{)}$ is a weight function.
Equation (LABEL:eq:dpo_gradient) highlights that DPO, while adaptively adjusting sample weights, inherently aims to increase the generation probability for chosen responses and decrease it for rejected ones. This approach can lead to two types of problems: 1) diminished generation probability for high-quality responses labeled as rejected, and 2) increased generation probability for low-quality responses labeled as chosen.
@@
Conversely, the more critical issue arises when low-quality responses are labeled as chosen. In such cases, their generation probabilities increase. This increase is particularly problematic because the probabilities of potential responses sum to one, meaning an increase in the probability of low-quality responses invariably decreases the share of high-quality responses. This shift substantially directs the learning process toward suboptimal outputs and declines the overall performance of LMs. A more detailed analysis of the sensitivity comparison between chosen and rejected responses will be provided in Appendix A.1.
Building upon these insights, fDPO uniquely addresses the issue of increased generation probability for low-quality chosen responses. It tackles these bottlenecks by discarding samples where the chosen responses are of lower quality compared to those generated by the LM $\pi_{\theta}$, as evaluated according to an RM. Through this process of consistent refinement, fDPO performs DPO on the improved dataset, thereby enhancing DPO’s effectiveness and ensuring a more effective alignment with human preferences.
@@ 5 Experiments
We first detail our experimental setup, including the pretrained models in Section 5.1 and the data in Section 5.2. We then evaluate the impact of data quality on DPO in Section 5.3 and the effectiveness of fDPO in Sections 5.4 and 5.5 in the context of instruction following tasks, focusing on the ability to generate appropriate responses to prompts. For our baseline comparison, we include a DPO implementation from the Transformer Reinforcement Learning (TRL) library. [^2] https://github.com/huggingface/trl

### v3

@@ Abstract
Filtered Direct Preference Optimization
Tetsuro Morimurathanks: Equal Contribution. Correspondence to: Tetsuro Morimura <morimura_tetsuro@cyberagent.co.jp>, Mitsuki Sakamoto <sakamoto_mitsuki@cyberagent.co.jp>,  Mitsuki Sakamoto11footnotemark: 1,   Yuu Jinnai,   Kenshi Abe,   Kaito Ariu CyberAgent
@@ 1 Introduction
Figure 1: Performance comparison of alignment methods using a 160M LM with the AlpacaFarm dataset (Dubois et al., 2023), where the gold rewards are adjusted so that the average reward of the initial LM is zero. (A) shows the impact of dataset quality on RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), with DPO exhibiting greater sensitivity to dataset quality variations. (B) compares the performance of DPO and the proposed fDPO on a mixed-quality dataset, illustrating that fDPO effectively mitigates the impact of data quality variations.
RLHF operates by taking a preference dataset and a language model (LM) as inputs to produce an LM refined by these preferences (Ouyang et al., 2022). It is broadly divided into two approaches concerning the use of a reward model (RM): RM-based RLHF, which learns an RM from the preference dataset and then uses it to optimize an LM through reinforcement learning (RL), and an RM-free approach that directly adjusts an LM based on preference data. This division mirrors the distinction between offline model-based and model-free RL (Sutton and Barto, 2018). [^1] RM-based RLHF first estimates the environment (specifically, the reward function; we do not need to estimate a state transition function because it is known in NLG tasks) and then optimizes an LM under the estimated environment. This approach is in itself a form of model-based RL. Each approach offers unique advantages and requires careful application based on specific goals and contexts. For instance, in scenarios with limited data, model-based RL might be preferable due to its data efficiency, though its computational cost is generally higher than that of model-free RL (Moerland et al., 2022; Levine et al., 2020). Consequently, RM-based RLHF may be more effective in leveraging data than RM-free methods, despite the higher computational cost and algorithmic complexity.
In this paper, we explore the impact of preference dataset quality on the performance of LMs optimized by DPO, specifically focusing on the quality of response texts rather than labeling accuracy. We demonstrate that DPO is more affected by text quality variations within the dataset than typical RLHF methods, as shown in Figure 1 (A). Notably, we observe that lower-quality data can create performance bottlenecks. In realistic applications of LLM alignment, the quality of responses can be highly diverse due to several factors such as differing skill levels among experts creating responses and the need to combine manually generated responses with those automatically generated by LLMs to manage annotation costs. This quality variation in response quality can severely impact performance of DPO.
In response to this challenge, we introduce a novel approach named filtered direct preference optimization (fDPO), which aims to harness potential data efficiency advantages of RM-based RLHF. It uses a trained RM to identify and discard samples of lower quality than those generated by an LM during fine-tuning. Our experiments show that fDPO significantly enhances the effectiveness of DPO, as illustrated in Figure 1 (B).
@@
  • Our experiments with two distinct datasets demonstrate that fDPO substantially enhances the performance of LMs.
The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 explains the background. In Section 4, we detail the proposed method, fDPO, explaining its mechanisms and the rationale behind its design. Section 5 presents the experimental results, illustrating the effectiveness of fDPO and its impact on LM performance. Finally, Section 6 concludes the paper, and Section 7 discusses limitations and directions for future work.
@@ 2 Related work
We examine methods for aligning LMs with human preferences, focusing on RLHF and its alternatives. Most RLHF approaches utilize an RM (Ouyang et al., 2022; Touvron et al., 2023; Dubois et al., 2023; Casper et al., 2023). These methods fine-tune LMs using RL algorithms such as REINFORCE (Williams, 1992; Rennie et al., 2017), proximal policy optimization (PPO) (Schulman et al., 2017), or their variants Sutton and Barto (2018). However, there are notable reinforcement-learning-free approaches (Zhao et al., 2023; Liu et al., 2024), and learning-free methods that leverage the RM at decoding time, with best-of-N (BoN) sampling being a prominent example (Stiennon et al., 2020; Nakano et al., 2021).
@@
DPO and its extensions (Azar et al., 2023; Tang et al., 2024; Pal et al., 2024; Singh et al., 2024) represent significant RM-free methods. Some DPO variants explore different regularizations (Wang et al., 2024) or use a divided dataset for stepwise training (Gou and Nguyen, 2024). Other variants propose adapting DPO online (Xu et al., 2023; Guo et al., 2024) or evaluating the quality difference between chosen and rejected responses for adding an offset to the DPO objective function (Amini et al., 2024) or incorporating curriculum learning (Gou and Nguyen, 2024). These approaches focus on response quality, which is relevant to our method.
Despite various advancements in DPO, the dependence on preference dataset quality has not been thoroughly analyzed. Our study aims to explore this significant dependence and attempts to refine the dataset for better performance. Additionally, our proposed fDPO method complements most of these developments. Integrating fDPO with these methods is an exciting possibility for future work, potentially leading to even more effective ways to align LMs with human preferences.
@@ 3 Background
This section explains RLHF in Section 3.1 and explores DPO in Section 3.2.
@@ 3.1 Reinforcement Learning from Human Feedback
Reinforcement learning from human feedback (RLHF) frames the application of human feedback to enhance performance of a language model (LM) within the context of an RL problem. The process incorporates a pre-trained LM $\pi_{\theta}(y\,|\,x)$, with $\theta$ denoting model parameters, $x$ the prompt, and $y$ the associated response. It also includes a demonstration dataset $\mathcal{D}_{\rm demo}$ for initial supervised fine-tuning and a preference dataset $\mathcal{D}$ for further RL fine-tuning. The aim is to refine the LM $\pi_{\theta}$ with these datasets $\mathcal{D}_{\rm demo}$ and $\mathcal{D}$. We will present an overview of the widely studied RLHF pipeline (Ouyang et al., 2022), establishing the notations and concepts for understanding our contributions. The RLHF pipeline comprises three principal phases: (i) supervised fine-tuning, (ii) reward modeling, and (iii) RL fine-tuning.
@@
To estimate the probability that a given response is preferred over another, the RM $r_{\phi}$ utilizes the Bradley-Terry model (Bradley and Terry, 1952), which is formulated as:
$p_{\textrm{BT}}$ $(y_{c}\succ y_{r}\,|\,x,r_{\phi})=\sigma(r_{\phi}(x,y_{c})-r_{\phi}(x,y_{r})),$
where $\sigma(x)=\frac{1}{1+\exp(-x)}$ is the sigmoid function. The RM is trained by maximizing the following log-likelihood of the observed preferences in the dataset:
$L(\phi)=\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}[\log\sigma(r_{\phi}(x,y_{c})-r_{\phi}(x,y_{r}))]$ (1)
@@
The RL fine-tuning phase uses the learned RM $r_{\phi}$ to optimize the SFT model $\pi_{\theta}$. The goal is to enhance $\pi_{\theta}$ by maximizing the expected reward while maintaining closeness to the reference LM $\pi_{\rm ref}$, striking a balance that avoids large deviations from the pre-trained behavior. The SFT model before RL fine-tuning is often used as $\pi_{\rm ref}$. This is achieved through policy gradient methods like proximal policy optimization (PPO) (Schulman et al., 2017). The optimization problem is formalized as
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\bigg{[}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\,|\,x)}[r_{\phi}(x,y)]$
$\hskip 55.48286pt-\beta D_{\textrm{KL}}(\pi_{\theta}(\cdot\,|\,x),\pi_{\rm ref}(\cdot\,|\,x))\bigg{]},$ (2)
@@
$D_{\textrm{KL}}(p,q)=\mathbb{E}_{y\sim p}\left[\log\frac{p(y)}{q(y)}\right].$
@@ 3.2 Direct Preference Optimization
$L_{\rm DPO}(\theta)$
$=\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\!\bigg{[}\log\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}$ (3)
$\hskip 99.58464pt-\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}\bigg{)}\bigg{]},$
@@ 4 Filtered Direct Preference Optimization
In this section, we propose an approach called filtered direct preference optimization (fDPO), which refines the dataset used in DPO. The principle of fDPO is straightforward: it aims to discard lower-quality samples compared to those generated by the LM. This strategy is intuitively derived from observing that lower-quality data can create performance bottlenecks in DPO. First, we give an implementation of fDPO in Section 4.1. Then, we will elaborate on the motivation of fDPO by analyzing DPO’s behavior in Section 4.2.
@@
fDPO needs to assess the quality of responses for filtering. For this purpose, a straightforward approach is to use an RM. This incorporation of an RM diverges from the RM-free nature of the original DPO, aligning fDPO closer to RM-based RLHF approaches and making DPO more effective in leveraging data.
@@
1:LM $\pi_{\theta}$, RM $r_{\phi}$, demonstration data $\mathcal{D}_{\rm demo}$, preference data $\mathcal{D}_{\rm pref}$, and maximum epoch $M$.
@@
6:while $m<M$ do
@@ 4.2 Background and Motivation for fDPO
The motivation for developing fDPO stems from the observation that the quality of data in DPO significantly affects the performance of the resulting LM. More specifically, upon differentiating the objective function of DPO in Eq. (3), we obtain
$\nabla_{\theta}L_{\rm DPO}(\theta)$ (4)
$=\beta\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\bigg{[}\underbrace{w_{\theta}(x,y_{c},y_{r})\nabla_{\theta}\log\pi_{\theta}(y_{c}|x)}_{\textrm{increase likelihood of }y_{c}}$
$\hskip 76.82243pt\underbrace{-w_{\theta}(x,y_{c},y_{r})\nabla_{\theta}\log\pi_{\theta}(y_{r}|x)}_{\textrm{decrease likelihood of }y_{r}}\bigg{]},$
where $w_{\theta}$ is a weight function defined as follows:
$w_{\theta}(x,y_{c},y_{r})$
$=\sigma\!\bigg{(}\!\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}-\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}\!\bigg{)}.$
Equation (4) highlights that DPO, while adaptively adjusting sample weights, inherently aims to increase the generation probability for chosen responses and decrease it for rejected ones. This approach can lead to two types of problems: 1) diminished generation probability for high-quality responses labeled as rejected, and 2) increased generation probability for low-quality responses labeled as chosen.
@@
Conversely, the more critical issue arises when low-quality responses are labeled as chosen. In such cases, their generation probabilities increase. This increase is particularly problematic because the probabilities of potential responses sum to one, meaning an increase in the probability of low-quality responses invariably decreases the share of high-quality responses. This shift substantially directs the learning process toward suboptimal outputs and declines the overall performance of LMs. A more detailed analysis of the sensitivity comparison between chosen and rejected responses will be provided in Appendix B.
Building upon these insights, fDPO effectively addresses the issue of increased generation probability for low-quality chosen responses. It tackles these bottlenecks by discarding samples where the chosen responses are of lower quality compared to those generated by the LM $\pi_{\theta}$, as evaluated according to an RM. Through this process of consistent refinement, fDPO performs DPO on the improved dataset, thereby enhancing DPO’s effectiveness and ensuring a more effective alignment with human preferences.
@@ 5 Experiments
We first detail our setup regarding pretrained models in Section 5.1 and datasets in Section 5.2. We then evaluate the impact of data quality on DPO in Section 5.3 and the effectiveness of fDPO in Section 5.4 on instruction following tasks using the AlpacaFarm dataset Dubois et al. (2023), focusing on the general ability to generate appropriate responses to prompts. Furthermore, we assess fDPO on the Anthropic HH datasets Bai et al. (2022) in Section 5.5, under a realistic setting where there are two types of responses: dataset responses and those generated by the SFT model. This setup closely mimics real-world applications, where the system must handle both pre-existing and newly generated responses. For our baseline comparison, we use DPO and PPO-based RLFH implementations from the Transformer Reinforcement Learning (TRL) library. [^2] https://github.com/huggingface/trl All experiments are conducted using a single NVIDIA A100 accelerator. The experiments using the AlpacaFarm dataset took approximately 3 days, and those using the Anthropic HH datasets took about 9 days of computation time. Details of the experimental parameters are provided in Appendix C.1.